40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Apr 2 04:11:20 2015, ericq, Update, LSC, Not much locking, Xover measurement CARMxOver_Apr3.pngApr2_Xover.xml.zip
    Reply  Thu Apr 2 15:34:34 2015, ericq, Update, LSC, Not much locking, Xover measurement xOverModel.pngloopModel.png
       Reply  Thu Apr 2 17:11:28 2015, ericq, Update, LSC, Not much locking, Xover measurement xOverModel.pngloopModel.png
Message ID: 11195     Entry time: Thu Apr 2 15:34:34 2015     In reply to: 11194     Reply to this: 11196
Author: ericq 
Type: Update 
Category: LSC 
Subject: Not much locking, Xover measurement 

Here's the comparison of last night's crossover measurement to my loop model. Not stellar, but not totally off base. All of the digital filters are read directly from the foton filter file, and translated from their SOS coefficients, so they should be accurate. I may have tallied together the wrong arrangement of FMs, though. I will recheck. 

Although I don't have a measurement to compare it with yet (as I don't know where the crossover was, the filter statesolder, etc. for the older loop measurements), here's what my current CARM loop model looks like, just for kicks. Here, only the first CM board boost is on. If we turn on some super boosting, we can probably ease up on some of the digital boosts, lower the crossover frequency, and put some lowpass that suppress the violin filters' effect on the crossover and reduces digital sensing noise injection. 

Lastly, I'll just note that my current MIST model predicts that the CARM cavity pole should be at ~170Hz, and a peak arm transmission of 180 times single arm power. I saw powers of ~120 last night. 

Attachment 1: xOverModel.png  210 kB  | Hide | Hide all
xOverModel.png
Attachment 2: loopModel.png  303 kB  | Hide | Hide all
loopModel.png
ELOG V3.1.3-